polling company reported that

33​%

of

2302

surveyed adults said that they play

basketball.basketball.

Complete parts​ (a) through​ (d) below.

a. What is the exact value that is
33​%

of

2302​?



Could the result from part​ (a) be the actual number of adults who said that they play

basketball question markbasketball?

Why or why​ not?
Answer #1

(a.) What is the exact value that is 33​% of 2302​?

required answer=2302*33/100=759.66

(b.) No,  the result from part​ (a) is not the actual number of adults who said that they play basketball.

since polling company reported that 33% of 2302 surveyed people.Here 2302 is sample size and 0.33 is the sample proportion. sample proportion would be the estimate of population proportion and if N is the population size then the actual number of population who said they play basketball=N*0.33